OpenStudy (destiny):

Marcus has 2 sisters.The sum of the girls' ages is 13.The product of their ages is 36.How old is each sister?

OpenStudy (anonymous):

Let the age of the sisters as a and b. then, a+b=13 and a*b=36 this implies a=13-b. hence (13-b)(b)=36. solve it then u get the answer

OpenStudy (anonymous):

Say one sister is Z and another sister is Y. You'd have to make a simultaneous equation, so it would be: Z+Y=13 ZY=36 You could solve it multiple ways, but the way I think is easiest would be substitution. Basically, if you used substitution, you would get Y=13-Z (because you solved for Y). Then you'd take that answer (13-Z) and substitute it for Y in the second equation. You'd then get: Z times (13-Z) = 36. So then you would just solve.

OpenStudy (destiny):

still no clue ....

OpenStudy (anonymous):

Z*(13-Z)=13Z-Z^2. 13Z-Z^2=36. Then you can factor I think. So: -Z^2+13Z-36=0 But we don't want a negative Z^2, so we have to flip all of the signs. Now we get: Z^2-13Z+36=0. Then factor. Because 13Z is negative and 36 is positive, you're going to have two negative answers in your factors. What two numbers multiply to 36 and add up to -13? -9 and -4. So Z can either equal 9 or 4, because you have to flip the signs. (And that should be apparent because you can't be a negative age.) So the sisters are aged 9 and 4. I hope that makes sense.. I haven't done that kind of equation in forever, but I think it's right.

OpenStudy (destiny):

thy...me too this is for my daughter she is in 4 g and I'm out of school for over 30 years..lol

OpenStudy (anonymous):

yes.sarahlynn is right. the answer should be 4 and 9 years old.

OpenStudy (destiny):

thank you this is the Destiny so one is 9 and the other one is 4 because 9*4=36 and 4*9=36 right so this would be right!!!